OpenStudy (anonymous):

1. A 500 mL sample of oxygen gas at 87.0°C and 500mmHg pressue is changed to new conditions at 447°C and 600 mmHg. What is the new volume? 2. A sample of helium gas has a volume of 100 mL at 27.0°C. If the pressure remains constant, what is the new volume when the temperature is raised to 127°C? 3. The temperature of a gas in a rigid container is increased from 25.0°C to 50.0°C. If the initial pressure is 101.3 kPa, what is the final pressure in kPa?

OpenStudy (anonymous):

@Abhisar Think you can handle these? ;) :P

OpenStudy (abhisar):

Hello @KissesFromKristine ! u can use this formula for finding the new formula \[\frac{ P1V1 }{ T1 } = \frac{ P2V2 }{ T2 }\]

OpenStudy (abhisar):

\[ \frac{ V1 }{ V2 } = \frac{ T1 }{ T2 }\] for finding the new volume when pressure is constant

OpenStudy (eric_d):

3)202.6?

OpenStudy (abhisar):

\[\frac{ P1 }{ P2 } = \frac{ T1 }{ T2 }\] for the third one

OpenStudy (anonymous):

I got 202.2, Eric. That correct @Abhisar ?

OpenStudy (abhisar):

Yes its correct answer for the third one

OpenStudy (abhisar):

Hav u figured out the rest two ? @KissesFromKristine

OpenStudy (anonymous):

Yay! Good job, @eric_d. We got it right!

OpenStudy (anonymous):

@Abhisar, no I haven't yet. :(

OpenStudy (abhisar):

why ? i gave u the formulas...just put the values and find out

OpenStudy (abhisar):

u there ?? @KissesFromKristine

OpenStudy (eric_d):

2) 470.37 ?

OpenStudy (anonymous):

Yeah, sorry. I'm still kind of struggling with aftermath of my sickness so I was taking some more medicine. I'll try them right now.

OpenStudy (abhisar):

okay..thats good :)

OpenStudy (abhisar):

Try them first urself

OpenStudy (abhisar):

okay..answer for the 3rd one will be 2140 ml

OpenStudy (anonymous):

for number 3? or the one we haven't done yet?

OpenStudy (abhisar):

sorry i meant number 1

OpenStudy (abhisar):

for Q.1 Ans=2140 Q2. Ans = 470 Q3. Ans= 202.2

OpenStudy (anonymous):

I agree with you on q2 & q3 but I don't know how you got 2140 for q1..

OpenStudy (abhisar):

okay..what are u getting for question no 1?

OpenStudy (anonymous):

Is it 1200 mL?

OpenStudy (abhisar):

how are u getting this ?

OpenStudy (abhisar):

are u using the formula i said ?

OpenStudy (anonymous):

I don't really understand the formula you said but this question on my sheet is multiple choice (not all of them are) and 2140 is not one of the answers. The possibilities are: A. 208 mL B. 300 mL C. 470 mL D. 833 mL E. 1.2 x 10^3 mL

OpenStudy (abhisar):

@ParthKohli

OpenStudy (anonymous):

Who is that?

OpenStudy (abhisar):

@Somy

OpenStudy (abhisar):

The first one @Somy

OpenStudy (somy):

which q? @Abhisar

OpenStudy (anonymous):

@Somy we're trying to figure out question 1

OpenStudy (somy):

hmmm let me check out

OpenStudy (abhisar):

ohkay

OpenStudy (somy):

its D

OpenStudy (somy):

try converting everything to standard units

OpenStudy (anonymous):

Ohh, i think I see what I did wrong. Hmm.. Thanks for the help @Somy & @Abhisar. I'm gonna post/try these last few questions then try to finally get some sleep.

OpenStudy (somy):

convert mL to dm3 convert mmHg to Pa and convert Celsius to Kelvin by adding 273 to the Celsius then put those values into the formula @Abhisar gave you :)

OpenStudy (abhisar):

and it will give me the answer ? @Somy

OpenStudy (somy):

yes :)

OpenStudy (eric_d):

1200

OpenStudy (abhisar):

How @eric_d ?

OpenStudy (abhisar):

It will be D as @Somy said....we just need to convert temperature from celsius into kelvin

OpenStudy (abhisar):

we can keep the rest as they are given...just celsius needs to be converted into kelvin

OpenStudy (anonymous):

D is what I got when I used Somy's method also. :)

OpenStudy (somy):

always when solving such questions make sure you use SI units to be on the safest side only then you can convert the unit of the result to your needed unit like here if u use dm3 in the formula your resulting volume will be in dm3 but u need cm3 so just multiply by 1000 and you'll get the answer :)

OpenStudy (abhisar):

u r ryt @Somy but we can avoid this conversion untill we r just comparing things and not finding absolute results. like in this question we were just comparing two equations. But we need to convert celsius into kelvin because 273.15 needs to be added and not multiplied both sides

OpenStudy (somy):

yes thats also right :)

OpenStudy (abhisar):

so in this question we will get the same results if we just convert temp into kelvin and keep the rest as they are

OpenStudy (abhisar):

Btw thankyou so much for bothering and clearing the doubt @Somy

OpenStudy (somy):

no problem :D @Abhisar